How much of a 35% solution should be mixed with 12ml of a 60% solution in order to become a 40% solution?
Answer by mananth(16946)

-------------- percent ---------------- Amount
Solution I 60 ---------------- 12 ml
Solution II 35 ---------------- x ml
Mixture 40 ---------------- 12 + x ml

60*12+35x=40(12 +x)
720+35x =480+40x
35x-40x =-720 + 480
-5 x = -240
/ -5
x = 48 ml Solution II

m.ananth@hotmail.ca